# Unsupervised Anomaly Detection through Mass Repulsing Optimal Transport

### v1

@@ 4.4 Robustness
Our method has two hyper-parameters, the entropic regularization penalty $\epsilon$, and the number of nearest neighbors $k$. In our experiments, we evaluated our methods over the values $\epsilon\in\{0,10^{-2},10^{-1},10^{0}\}$, where $\epsilon=0$ implies using exact OT, i.e., linear programming. For MROT, we use $k\in\{5,10,20,\cdots,50\}$. Note that, while $\epsilon$ is linked to the transportation plan, $k$ is linked to the ground-cost. We summarize our results in Figure 9, where we present KDEs over the AUC-ROC scores of each dataset, for each combination of hyper-parameters.
@@ 5 Conclusion
In this paper, we introduced a novel, general purpose anomaly detection algorithm based on optimal transport theory. Our method works under the assumption that the local neighborhood of anomalous samples is likely more irregular than that of normal samples. Based on this idea, we engineer the ground-cost in optimal transport, to encourage samples to send their mass just outside an exclusion zone, defined through its $k-$nearest neighbors. While this idea bears some similarity to optimal transport with repulsive costs (Di Marino et al., 2017), the defined ground-cost is not repulsive, and, as we show in our experiments, it leads to better anomaly detectors. We thoroughly experiment on the AdBench (Han et al., 2022) benchmark, and the Tennessee Eastman process (Downs and Vogel, 1993; Reinartz et al., 2021; Montesuma et al., 2024b), showing that our method outperforms previously proposed methods. However, as we highlight in section 4.2, our method inherits the limitations of classic optimal transport (Montesuma et al., 2024a). Especially, it needs $\mathcal{O}(n^{2})$ computational storage, has $\mathcal{O}(n^{3}\log n)$ computational complexity and is difficult to estimate in high dimensions. Nonetheless, as our experiments demonstrate, our method is practical and performative in real-world anomaly detection scenarios.
@@ Appendix A Additional Details on Experiments
A.1 Additional results on AdBench

### v2

@@ 4.3 Ablations
We give an additional reasoning on the choice of $\epsilon$. From the classical OT (Peyré and Cuturi, 2020, Chapter 4), using $\epsilon>0$ leads to a nonsparse plan. This means that there are more indices $(i,j)$, for which $\gamma_{ij}>0$. Likewise, if we let $\epsilon\rightarrow+\infty$, then $\gamma$ converges to the trivial coupling $\gamma_{ij}=p_{i}p_{j}$. Assuming uniform importance, i.e., $p_{i}=n^{-1}$, and plugging these results back into equation 10, we have $t_{i}=n^{-1}\sum_{j=1}^{n}\tilde{c}_{k}(x_{i}^{(P)},x_{j}^{(P)})$, that is, for a large entropic penalty, we end up simply computing the average cost of each sample. The implication of this analysis is as follows: the MROT plan captures some structure in the relationship between anomalous and normal points, which contributes to a more accurate anomaly score.
Robustness to regression algorithm. As we mentioned in our toy example, it is necessary to use a nonlinear regression model for generalizing the anomaly score, as the relationship between $\mathbf{x}_{i}^{(P)}$ and $\hat{F}_{\mathcal{T}}(t_{i})$ is likely nonlinear. We ablate on the choice of regression algorithm, exploring the use of gradient boosting (Friedman, 2002), SVR (Smola and Schölkopf, 2004) and k-nearest neighbors regression. We show a summary of our results in Figure 13. Note that, in general, the performance of boosting and SVR is similar and stable among different choices for $k$ in $\mathcal{N}_{k}$ in our engineered cost. The performance of $k-$nearest neighbors regression is sub-optimal, especially due to overfitting.
Figure 13: Ablation of MROT performance with respect regression model choice. Overall, gradient boosting and SVR have similar performance, whereas $k-$nearest neighbors has sub-optimal performance due to overfitting.
In addition to the pipeline shown in Algorithm 1, it is possible to add a model selection step, e.g., through k-fold cross-validation. The idea is to partition the training dataset $\{\mathbf{x}_{i}^{(P)},\hat{F}_{\mathcal{T}}(t_{i})\}_{i=1}^{n}$ into a training and a validation set. Then, among a pool of possible regression models, the one that has the minimum error in the validation set.
@@ 5 Conclusion
In this paper, we introduced a novel, general purpose anomaly detection algorithm based on optimal transport theory. Our method works under the assumption that the local neighborhood of anomalous samples is likely more irregular than that of normal samples. Based on this idea, we engineer the ground-cost in optimal transport, to encourage samples to send their mass just outside an exclusion zone, defined through its $k-$nearest neighbors. While this idea bears some similarity to optimal transport with repulsive costs (Di Marino et al., 2017), the defined ground-cost is not repulsive, and, as we show in our experiments, it leads to better anomaly detectors. We thoroughly experiment on the AdBench (Han et al., 2022) benchmark, and the Tennessee Eastman process (Downs and Vogel, 1993; Reinartz et al., 2021; Montesuma et al., 2024b), showing that our method outperforms previously proposed methods in real-world benchmarks.
Limitations. Our methods faces 2 limitations from OT computation. First, we need to store a $n\times m$ matrix, which incurs a $\mathcal{O}(n^{2})$ storage complexity. Furthermore, computing $\gamma_{\epsilon}^{\star}$ has $\mathcal{O}(n^{3}\log n)$ computational complexity. For that reason, in the AdBench experiments, we had to downsample larger datasets to $n=20,000$ samples. Second, OT is notoriously hard to estimate in higher dimensions. As a result, our method struggles with high-dimensional features, such as those in the NLP datasets of the AdBench benchmark.
In the literature, there are different ways of tackling these limitations, but we leave those for future works. For instance, a workaround to the first limitations is to focus on mini-batch OT (Fatras et al., 2020) or computing OT with neural nets (Makkuva et al., 2020; Korotin et al., 2023). Likewise, the computation of OT in high dimensions can be done, for instance, through modifications of the original OT objective, such as subspace robust OT (Paty and Cuturi, 2019).
@@ Appendix A Additional Details on Experiments
A.1 Detailed results on AdBench